OpenStudy (anonymous):

if x=3, then x square =9 A) x square =9 if and only if x=3, B) if x square, then x=3, C) x=3 if and only if x square =9, D) This statement is not reversible

OpenStudy (texaschic101):

(A) not true....because x could = -3 and x^2 would still = 9 (B) not always true....if x^2 = 9, then x could = - 3 (C) Not always true....x could = -3 and x^2 would be 9 That leaves D. You might want to get a second opinion.

OpenStudy (texaschic101):

@cwrw238 ...can you take a peek at this, please

OpenStudy (texaschic101):

@tcarroll010 .....please look at this

OpenStudy (texaschic101):

@jim_thompson5910 ...will you help

OpenStudy (texaschic101):

I suppose it could also be B. Using the converse statement.